# Judging LLMs on a Simplex

### v1

@@ D.3 LLM candidates
For consistent comparison across datasets, we evaluated the set of candidate LLMs shown in Table D.2.
Table D.2: LLM candidates evaluated across all experimental settings
@@
For the MTBench dataset [50], we assess the candidates in the provided dataset: GPT-4-0613, Claude-1, Llama-2-13B-Chat, Vicuna-13B, and Alpaca-13B. Similarly, for the TLDR (SummEval) benchmark [11], we assess the 12 provided language models.
@@ D.3.1 Comparison Methods
Simple Averaging with Bootstrap Confidence Intervals: This comparator uses the bootstrap approach for population ranks proposed in [48] to generate confidence intervals for the simple averaging estimate.
Pairwise Comparison Approach: As a representative pairwise comparison method, we an extension of the Bradley-Terry model with ties [31], which estimates candidate ability parameters based on win-loss-tie patterns in pairwise evaluations. We calculate confidence intervals for this approach using bootstrap.
@@ D.4.2 Human Judgement Prompts
For MTBench, we use a prompt similar to the original prompt given to human judges as described in [50] for the single-turn and two-turn cases. Similar to the binary tasks, we replace the [[question]] and [[candidates_section]] placeholders with the prompt and the conversation history using the following format:
@@
For the TLDR (SummEval) dataset, we use the same evaluation framework as in [11] that assesses news article summarization quality across four distinct criteria: relevance, consistency, fluency, and coherence. Each dimension is evaluated on a 5-point Likert scale with specific definitions provided to ensure consistent interpretation:
@@
* Fluency: This rating measures the quality of individual sentences-are they well-written and grammatically correct?
* Coherence: This rating measures the quality of all sentences collectively-do they fit together and sound natural? Consider the quality of the summary as a whole.
@@
// Add evaluations for Candidate #2, #3, ... up to #[[num_candidates]] following the same structure
@@ D.5 MTBench self-preference
Table D.3 presents the frequency of scores (ranging from 1, low, to 5, high) assigned by the two LLM judges—Claude 3.5 Haiku and GPT-4o mini—to various candidate LLMs based on their responses to two-turn questions from the MTBench dataset. The results suggest potential self-preference biases in the LLM-based evaluations. The Claude 3.5 Haiku judge awarded its predecessor, Claude v1, a high frequency of top scores (48 instances of ’5’), surpassing other models like GPT-4 (40 instances of ’5’). The GPT-4o mini judge assigned an equal number of perfect ’5’ scores (50 instances each) to both its own family model, GPT-4, and to Claude v1. While GPT-4o mini gives the same number of top scores to GPT-4 and Claude, the scoores from the Claude judge suggests that models tend to receive more favorable evaluations from judges within the same model family, highlighting a self-preference.
Table D.3: Distribution of scores (1-5) assigned by Claude 3.5 Haiku and GPT-4o mini judges to candidate LLMs on the MTBench two-turn benchmark. Cell values represent the frequency of each score.
@@ D.6 Score Mapping for Simplex Visualizations
$f_{\text{MTBench}}(s)=\begin{cases}1,&\text{if }s\in\{1,2\}\\ 2,&\text{if }s\in\{0,3\}\\ 3,&\text{if }s\in\{4,5\}\\ \end{cases}$
@@
Appendix E Proofs
@@ E.1 Proof for Theorem 1
E.2 Proof for Theorem 2
@@
$\left(\Theta^{-1}(\gamma_{1}-\gamma_{2})\right)^{\top}\left(\begin{matrix}0\\ 1\\ 2\end{matrix}\right)=0.$ (E.4)
@@
$\left(\left(\Theta+h\Delta\right)^{-1}(\gamma_{1}-\gamma_{2})\right)^{\top}\left(\begin{matrix}0\\ 1\\ 2\end{matrix}\right).$ (E.5)
To prove nonidentifiability of the rankings, it thus suffices to show that for any $\Theta$, there exists $\gamma_{1},\gamma_{2},\Delta$ such that the score difference (E.5) is zero at $h=0$, i.e. (E.4) holds, and the derivative of the score difference (E.5) is nonzero, i.e.
$\nabla_{h}\left(\left(\Theta+h\Delta\right)^{-1}(\gamma_{1}-\gamma_{2})\right)^{\top}\left(\begin{matrix}0\\ 1\\ 2\end{matrix}\right)\propto\left(\Theta^{-1}(\gamma_{1}-\gamma_{2})\right)^{\top}(\Theta^{-1}\Delta)^{\top}\left(\begin{matrix}0\\ 1\\ 2\end{matrix}\right)\neq 0.$ (E.6)
@@
Let $\gamma_{k}=\Theta\pi_{k}$ for $k=1,2$. Then (E.5)=0italic-(E.5italic-)$=0$ at $h=0$. To prove that (E.6)0italic-(E.6italic-)$\neq 0$ at $h=0$, let $u=(\Theta^{-1})^{T}\left(\begin{matrix}0\\ 1\\ 2\end{matrix}\right)$ and $\bar{u}=\frac{\mathbf{1}^{T}u}{3}$. Define
@@
[…]
@@
E.3 Proof for asymptotic consistency of independent model
@@
$\prod_{i=1}^{n}\prod_{j=1}^{J}\prod_{k\neq j}\Big{[}\sum_{m=1}^{M}\underbrace{\Pr\left(\hat{S}^{(j)}_{ik}|S_{ik}^{*}=m;\theta_{m,k}^{(j)}\right)}_{=\theta_{m,k,\hat{S}^{(j)}_{ik}}^{(j)}}\underbrace{\Pr(S_{ik}^{*}=m)}_{=\pi_{k,m}}\Big{]}.$ (E.7)
Here we prove that the candidate ranking that maximizes (E.7) still converges asymptotically to the true ranking as long as the true rank is identifiable, despite the probability model ignoring correlation between judge-assigned scores (and despite the fact that $\theta$ and $\pi$ are not identifiable). This can be viewed as an extension of traditional statistical results that the MLE for parameters that only depend on the first-moment is asymptotically consistent even when we do not correctly model the second-moments or higher. This result is stated more formally below. For simplicity, we present the result for one judge that satisfies Assumption 1. Extensions to multiple judges and to use Assumption 1 rather than 1 are straightforward.
@@
Suppose Assumption 1 holds for one judge. The estimated ranking $\hat{\xi}_{n}$ using scores from the judge by maximizing (4.2) converges to the true ranking $\xi^{*}$ as $n\rightarrow\infty$ as long as:
@@
$\max_{\gamma_{1},\cdots,\gamma_{K}}\prod_{i=1}^{n}\prod_{k=1}^{K}\Pr\left(\hat{S}_{i,k};\gamma_{k}\right).$ (E.8)
@@
Because the solution to (E.8) corresponds to the MLE for a multinomial model, and standard regularity conditions apply, the estimator is asymptotically consistent. Formally, this implies that the estimated parameters $\hat{\gamma}_{k,n}$ satisfy
$\Pr\left(\ell_{n}(\gamma_{1},\cdots,\gamma_{K})-\ell_{n}(\gamma_{1}^{*},\cdots,\gamma_{K}^{*})<0\quad\forall\|\gamma-\gamma^{*}\|\geq a\right)\rightarrow 1$ (E.9)
@@
Let’s analyze the optimization problem (4.2) optimized with respect to the parameters $(\theta,\delta,\xi)$ where $\xi$ is the ranking of the candidate, $\delta\in\mathbb{R}^{K,+}$, and $\pi_{k}=\sum_{i=1}^{K}\mathbbm{1}\{i<=\xi_{k}\}\delta_{i}$. Let a function $g$ denote the mapping from a parameter set $(\theta,\delta,\xi)$ describing judge and candidates to the set of marginal parameters $\gamma_{1},\cdots,\gamma_{K}$. Under Assumption 1, there is some true set of parameters $(\theta^{*},\delta^{*},\xi^{*})$ with corresponding marginal parameters $\gamma_{1}^{*},\cdots,\gamma_{K}^{*}$. Then because the set of models considered in (4.2) is a subset of that in (E.8), we have that
$\Pr\left(\ell_{n}(g(\theta,\delta,\xi))-\ell_{n}(g(\theta^{*},\delta^{*},\xi^{*}))<0\quad\forall\left\|g(\theta,\delta,\xi)-q(\theta^{*},\delta^{*},\xi^{*})\right\|\geq a\right)$ (E.10)
$\geq$ $\Pr\left(\ell_{n}(\gamma_{1},\cdots,\gamma_{K})-\ell_{n}(\gamma_{1}^{*},\cdots,\gamma_{K}^{*})<0\quad\forall\|\gamma-\gamma^{*}\|\geq a\right).$ (E.11)
Because (E.11) converges to one for any $a>0$, then (E.10) converges asymptotically to one as well for all $a>0$. Taking $a\rightarrow 0$, we have that
$\sup_{\theta_{n},\delta_{n},\xi_{n}}g(\hat{\theta}_{n},\delta_{n},\xi_{n})-(\gamma_{1}^{*},\cdots,\gamma_{K}^{*})\rightarrow_{p}0.$ (E.12)
Furthermore, because we assumed that (i) the true ranking $\xi^{*}$ is identifiable from the marginals and (ii) no two candidates have the same exact true score, then the mapping from candidate marginals $(\gamma_{1},\cdots,\gamma_{K})$ to candidate rankings is unique and locally constant at $(\gamma_{1}^{*},\cdots,\gamma_{K}^{*})$. Then by the continuous mapping theorem, we have that the estimated rankings $\hat{\xi}_{n}$ when maximizing (4.2) converges asymptotically to the true rankings $\xi^{*}$. ∎

### v2

@@ D.3 LLM candidates
For consistent comparison across datasets, we evaluated the set of candidate LLMs shown in Table 2.
Table 2: LLM candidates evaluated across all experimental settings
@@
For the MTBench dataset (Zheng2023-wr), we assess the candidates in the provided dataset: GPT-4-0613, Claude-1, Llama-2-13B-Chat, Vicuna-13B, and Alpaca-13B. Similarly, for the TLDR (SummEval) benchmark (Fabbri2021-ie), we assess the 12 provided language models.
@@ D.3.1 Comparison Methods
Simple Averaging with Bootstrap Confidence Intervals: This comparator uses the bootstrap approach for population ranks proposed in Xie2009-wv to generate confidence intervals for the simple averaging estimate.
Pairwise Comparison Approach: As a representative pairwise comparison method, we an extension of the Bradley-Terry model with ties (Rao1967-fv), which estimates candidate ability parameters based on win-loss-tie patterns in pairwise evaluations. We calculate confidence intervals for this approach using bootstrap.
Prediction-Powered Inference (PPI): For benchmarks with ground truth labels, we implement PPI (Angelopoulos2023-tv) as an alternative ranking method. PPI uses a small labeled dataset to calibrate predictions from LLM judges on a larger unlabeled dataset, providing statistically valid confidence intervals for candidate rankings. Our implementation: (i) randomly partitions questions into labeled and unlabeled sets (using 5% or 10% labeled fractions), (ii) applies the PPI mean estimator using the ppi_py library with judge scores converted to a 0-1 scale, and (iii) generates confidence intervals through bootstrap resampling (300 iterations) while maintaining the fixed labeled/unlabeled partition. We apply PPI to GPQA, MMLU Pro, and Omni-MATH benchmarks where ground truth answers are available. Results for different labeled data fractions are presented in Section E.1.
@@ D.4.2 Human Judgement Prompts
For MTBench, we use a prompt similar to the original prompt given to human judges as described in (Zheng2023-wr) for the single-turn and two-turn cases. Similar to the binary tasks, we replace the [[question]] and [[candidates_section]] placeholders with the prompt and the conversation history using the following format:
@@
For the TLDR (SummEval) dataset, we use the same evaluation framework as in (Fabbri2021-ie) that assesses news article summarization quality across four distinct criteria: relevance, consistency, fluency, and coherence. Each dimension is evaluated on a 5-point Likert scale with specific definitions provided to ensure consistent interpretation:
@@
* Fluency: This rating measures the quality of individual sentences: are they well-written and grammatically correct?
* Coherence: This rating measures the quality of all sentences collectively: do they fit together and sound natural? Consider the quality of the summary as a whole.
@@
// Add evaluations for Candidate #2, #3, up to #[[num_candidates]] following the same structure
@@ D.5 MTBench self-preference
Table 3 presents the frequency of scores (ranging from 1, low, to 5, high) assigned by the two LLM judges—Claude 3.5 Haiku and GPT-4o mini—to various candidate LLMs based on their responses to two-turn questions from the MTBench dataset. The results suggest potential self-preference biases in the LLM-based evaluations. The Claude 3.5 Haiku judge awarded its predecessor, Claude v1, a high frequency of top scores (48 instances of ’5’), surpassing other models like GPT-4 (40 instances of ’5’). The GPT-4o mini judge assigned an equal number of perfect ’5’ scores (50 instances each) to both its own family model, GPT-4, and to Claude v1. While GPT-4o mini gives the same number of top scores to GPT-4 and Claude, the scoores from the Claude judge suggests that models tend to receive more favorable evaluations from judges within the same model family, highlighting a self-preference.
Table 3: Distribution of scores (1-5) assigned by Claude 3.5 Haiku and GPT-4o mini judges to candidate LLMs on the MTBench two-turn benchmark. Cell values represent the frequency of each score.
@@ D.6 Score Mapping for Simplex Visualizations
[…]
@@
Appendix E Additional Results
E.1 Prediction-Powered Inference Results
Table 4 presents ranking results using Prediction-Powered Inference (PPI) across GPQA, MMLU Pro, and Omni-MATH benchmarks with varying amounts of labeled data (5% and 10%). PPI demonstrates strong performance, achieving high correlations with true rankings and consistently strong coverage rates. These results confirm that PPI is highly effective at quantifying uncertainty when calibrated with even small amounts of labeled data, providing reliable confidence intervals while maintaining competitive ranking accuracy. The method’s strength lies in its ability to leverage limited ground truth to correct for judge bias, making it an excellent choice when some labeled examples are available. In contrast, the Bayesian framework provides robust uncertainty quantification when gold standard labels are unavailable, addressing the complementary scenario where obtaining any ground truth is impractical or infeasible.
Table 4: Performance of PPI ranking method across benchmark datasets with different labeled fractions. Results show Spearman correlation (Corr) and coverage rates (Cov) with 95% confidence intervals.
Labeled Fraction GPQA MMLU Pro Omni-MATH Corr Cov Corr Cov Corr Cov 5% 0.773 (0.445, 0.889) 0.889 (0.778, 1.000) 0.902 (0.691, 0.919) 0.947 (0.749, 1.000) 0.789 (0.486, 0.905) 0.895 (0.789, 1.000) 10% 0.875 (0.529, 0.923) 0.944 (0.790, 1.000) 0.922 (0.782, 0.929) 1.000 (0.737, 0.947) 0.737 (0.621, 0.929) 0.895 (0.801, 1.000)
E.2 Posterior Distribution Plots
Figures E.11-E.13 present posterior distributions for candidate rankings across additional benchmark datasets, complementing the GPQA results shown in the main text. These visualizations reveal consistent patterns in how different methods quantify uncertainty.
MMLU Pro and Omni-MATH (FigureE.11): Both datasets exhibit similar characteristics to GPQA, with the Bayesian method producing notably wider credible intervals that capture epistemic uncertainty about judge quality. The increased width is particularly pronounced for middle-ranked candidates, where judge disagreement tends to be highest. For Omni-MATH specifically, we observe greater overall uncertainty across all methods, consistent with the sensitivity analysis showing this dataset’s vulnerability to constancy violations.
Figure E.11: Posterior distributions for candidate rankings on MMLU Pro (left) and Omni-MATH (right), with candidates ordered by true ranking (best at top).
SummEval (Figure E.13): Across the four evaluation dimensions (coherence, consistency, fluency, and relevance), the Bayesian approach maintains wider intervals while preserving accurate point estimates. The coherence and consistency dimensions show tighter clustering of posterior distributions compared to fluency and relevance, suggesting these aspects may be more reliably evaluated by LLM judges. The top-ranked candidates (M22, M23, M17) show relatively narrow intervals across all methods, indicating strong agreement on the best performers.
Figure E.12: Posterior distributions for candidate rankings on MTBench, with candidates ordered by true ranking (best at top).
MTBench (Figure E.12): The MTBench dataset exhibits tight posterior distributions across all methods, indicating strong judge consensus. Even the Bayesian approach produces relatively narrow credible intervals, suggesting low epistemic uncertainty for this dataset. This aligns with our sensitivity analysis showing MTBench maintains stable rankings under relaxed constancy assumptions, representing a case where judge-based evaluation proves particularly reliable.
A key pattern across all datasets is that while the three methods generally agree on median rankings, they diverge substantially in uncertainty quantification. The Bootstrap and Bradley-Terry methods produce similar narrow confidence intervals that account only for sampling variability. In contrast, the Bayesian framework’s wider intervals reflect both aleatoric and epistemic uncertainty, explaining its superior coverage rates reported in main text. This difference is most pronounced for candidates with intermediate performance levels, where judge quality uncertainty has the greatest impact on ranking uncertainty.
Figure E.13: Posterior distributions for candidate rankings on TLDR/SummEval across four evaluation dimensions: coherence (top left), consistency (top right), fluency (bottom left), and relevance (bottom right). Candidates are ordered by true ranking (best at top).
E.3 Robustness to Self-Preference Bias
While our main experiments demonstrate the framework’s ability to handle naturally occurring judge biases (e.g., MTBench self-preference in Table 3), we conducted additional controlled experiments to systematically evaluate robustness to self-preference bias.
Experimental Design: Using the GPQA dataset, we artificially induced self-preference bias by manipulating judge-assigned scores when judges evaluated candidates from the same model family. Specifically, when Claude judges evaluated Claude candidates or GPT judges evaluated GPT candidates, we upward-shifted the assigned score with probability 0.8. This creates a strong but realistic bias pattern similar to that observed in real evaluations.
Bias Injection Protocol:
  • For binary scores (correct/incorrect), incorrect answers were changed to correct with probability 0.8
  • For abstentions, we converted them to correct with probability 0.8
  • Non-self evaluations remained unchanged
This manipulation creates a challenging scenario where traditional averaging methods would systematically overestimate the performance of judge-candidates.
Table 5: Performance under simulated self-preference bias on GPQA dataset
Method Correlation Coverage
Bayesian with self-preference adjustment 0.811 0.852
Bayesian without adjustment 0.701 0.611
Bootstrap 0.680 0.519
Bradley-Terry 0.673 0.482
Simple Average 0.680 0.056
Single Judge 0.683 0.148
The results demonstrate that the Bayesian framework with self-preference adjustment effectively corrects for the induced bias, achieving correlation and coverage rates comparable to the unbiased setting shown in the main text. Without adjustment, all methods suffer substantial performance degradation, with the Bayesian method’s correlation dropping from 0.881 to 0.701.
Appendix F Proofs
@@ F.1 Proof for Theorem 1
F.2 Necessary and Sufficient Conditions for 2-Level Systems
While Theorem 1 provides sufficient conditions for ranking identifiability under constancy assumptions, we can characterize the exact necessary and sufficient conditions for when rankings are preserved in 2-level systems with a single judge.
Lemma 1.
Let $(k_{1},k_{2},\ldots,k_{K})$ be a permutation such that the true scores satisfy $\Pr(S_{k_{1}}^{*}=2)>\Pr(S_{k_{2}}^{*}=2)>\cdots>\Pr(S_{k_{K}}^{*}=2)$. Then the following are equivalent:
  1. (i)
    The judge-assigned scores preserve this ordering: $\Pr(\hat{S}_{k_{1}}^{(j)}=2)>\Pr(\hat{S}_{k_{2}}^{(j)}=2)>\cdots>\Pr(\hat{S}_{k_{K}}^{(j)}=2)$
  2. (ii)
    For all $\ell\in\{1,\ldots,K-1\}$: $\Pr(\hat{S}_{k_{\ell}}^{(j)}=2)>\Pr(\hat{S}_{k_{\ell+1}}^{(j)}=2)$
Proof.
The forward direction (i $\Rightarrow$ ii) is immediate. For (ii $\Rightarrow$ i), assume $Pr(\hat{S}_{k_{\ell}}=2)>Pr(\hat{S}_{k_{\ell+1}}=2)$ for every $\ell=1,\dots,K-1$. Fix any indices $1\leq i<j\leq K$. By repeated transitivity,
$\hat{p}_{k_{i}}>\hat{p}_{k_{i+1}}>\cdots>\hat{p}_{k_{j}},$
hence $\hat{p}_{k_{i}}>\hat{p}_{k_{j}}$. Since this holds for all $i<j$, we obtain the strict chain $\hat{p}_{k_{1}}>\hat{p}_{k_{2}}>\cdots>\hat{p}_{k_{K}}$, proving (i). ∎
The key insight is that we only need to verify K1K-1 local inequalities between adjacent candidates rather than all $\binom{K}{2}$ pairwise comparisons. For condition (ii) to hold given the true ranking, we require:
Theorem 4 (Necessary and Sufficient Conditions).
Consider a 2-level scoring system with a single judge. True rankings are preserved by judge $j$ for all possible candidate configurations if and only if for each adjacent pair $(k_{\ell},k_{\ell+1})$ in the true ranking:
$\theta_{2,k_{\ell}}^{(1)}>\theta_{1,k_{\ell+1}}^{(1)}\quad\text{(cross-term dominance)}$ (9)
$\theta_{1,k_{\ell}}^{(1)}\geq\theta_{1,k_{\ell+1}}^{(1)},\text{ and if equal, then }\theta_{2,k_{\ell}}^{(1)}>\theta_{1,k_{\ell}}^{(1)}$ (10)
$\theta_{2,k_{\ell}}^{(1)}\geq\theta_{2,k_{\ell+1}}^{(1)},\text{ and if equal, then }\theta_{2,k_{\ell+1}}^{(1)}>\theta_{1,k_{\ell+1}}^{(1)}$ (11)
where $\theta_{m,k}^{(1)}=\Pr(\hat{S}_{k}^{(1)}=2|S_{k}^{*}=m)$ represents the judge’s confusion matrix entries.
Proof.
Condition (ii) from Lemma 1 states that $\Pr(S_{k_{\ell}}^{*}=2)>\Pr(S_{k_{\ell+1}}^{*}=2)$ implies $\Pr(\hat{S}_{k_{\ell}}^{(1)}=2)>\Pr(\hat{S}_{k_{\ell+1}}^{(1)}=2)$. Using the notation $\pi_{k,m}=\Pr(S_{k}^{*}=m)$ and $\theta_{m,k}^{(1)}=\Pr(\hat{S}_{k}^{(1)}=2|S_{k}^{*}=m)$, this becomes:
$\pi_{k_{\ell},2}>\pi_{k_{\ell+1},2}\implies\gamma_{k_{\ell}}^{(1)}>\gamma_{k_{\ell+1}}^{(1)}$ (12)
where $\gamma_{k}^{(1)}=\theta_{1,k}^{(1)}(1-\pi_{k,2})+\theta_{2,k}^{(1)}\pi_{k,2}$ is the judge-assigned score probability.
For (12) to hold, we need the difference
$g(\pi_{k_{\ell},2},\pi_{k_{\ell+1},2})$ $=\gamma_{k_{\ell}}^{(1)}-\gamma_{k_{\ell+1}}^{(1)}$
$=\theta_{1,k_{\ell}}^{(1)}(1-\pi_{k_{\ell},2})+\theta_{2,k_{\ell}}^{(1)}\pi_{k_{\ell},2}-\theta_{1,k_{\ell+1}}^{(1)}(1-\pi_{k_{\ell+1},2})-\theta_{2,k_{\ell+1}}^{(1)}\pi_{k_{\ell+1},2}$
$=(\theta_{1,k_{\ell}}^{(1)}-\theta_{1,k_{\ell+1}}^{(1)})+(\theta_{2,k_{\ell}}^{(1)}-\theta_{1,k_{\ell}}^{(1)})\pi_{k_{\ell},2}+(\theta_{1,k_{\ell+1}}^{(1)}-\theta_{2,k_{\ell+1}}^{(1)})\pi_{k_{\ell+1},2}$ (13)
to be positive for all $\pi_{k_{\ell},2}>\pi_{k_{\ell+1},2}$ with $0\leq\pi_{k_{\ell+1},2}\leq\pi_{k_{\ell},2}\leq 1$.
Since $g(\pi_{k_{\ell},2},\pi_{k_{\ell+1},2})$ is linear in both arguments, it is positive throughout the feasible region if and only if it is non-negative at all extreme points. The feasible region $\{(\pi_{k_{\ell},2},\pi_{k_{\ell+1},2}):0\leq\pi_{k_{\ell+1},2}\leq\pi_{k_{\ell},2}\leq 1\}$ has three extreme points:
At $(0,0)$: $g(0,0)=\theta_{1,k_{\ell}}^{(1)}-\theta_{1,k_{\ell+1}}^{(1)}\geq 0$
At $(1,0)$: $g(1,0)=\theta_{1,k_{\ell}}^{(1)}-\theta_{1,k_{\ell+1}}^{(1)}+(\theta_{2,k_{\ell}}^{(1)}-\theta_{1,k_{\ell}}^{(1)})=\theta_{2,k_{\ell}}^{(1)}-\theta_{1,k_{\ell+1}}^{(1)}>0$
At $(1,1)$: $g(1,1)=\theta_{1,k_{\ell}}^{(1)}-\theta_{1,k_{\ell+1}}^{(1)}+(\theta_{2,k_{\ell}}^{(1)}-\theta_{1,k_{\ell}}^{(1)})+(\theta_{1,k_{\ell+1}}^{(1)}-\theta_{2,k_{\ell+1}}^{(1)})=\theta_{2,k_{\ell}}^{(1)}-\theta_{2,k_{\ell+1}}^{(1)}\geq 0$
Note that $g(1,0)$ must be strictly positive since this corresponds to the interior of the feasible region where $\pi_{k_{\ell},2}>\pi_{k_{\ell+1},2}$.
For strict inequality in (12) when $\pi_{k_{\ell},2}>\pi_{k_{\ell+1},2}$, we also need $g$ to be strictly positive when moving slightly away from the boundary points. This requires:
  • If $\theta_{1,k_{\ell}}^{(1)}=\theta_{1,k_{\ell+1}}^{(1)}$, then $\theta_{2,k_{\ell}}^{(1)}>\theta_{1,k_{\ell}}^{(1)}$ to ensure $g(\Delta,0)>0$ for small $\Delta>0$
  • If $\theta_{2,k_{\ell}}^{(1)}=\theta_{2,k_{\ell+1}}^{(1)}$, then $\theta_{2,k_{\ell+1}}^{(1)}>\theta_{1,k_{\ell+1}}^{(1)}$ to ensure $g(1,1-\Delta)>0$ for small $\Delta>0$
These conditions correspond exactly to equations (9), (10), and (11). ∎
Remark 1.
These necessary and sufficient conditions reveal that without constancy assumptions, ranking preservation requires complex coordination between confusion matrices: higher-ranked candidates must have both larger false-positive rates (FPR) and true-positive rates (TPR) than lower-ranked candidates, with the higher rank’s TPR exceeding the lower rank’s FPR.
F.3 Proof for Theorem 2
@@
$\left(\Theta^{-1}(\gamma_{1}-\gamma_{2})\right)^{\top}\left(\begin{matrix}0\\ 1\\ 2\end{matrix}\right)=0.$ (14)
@@
$\left(\left(\Theta+h\Delta\right)^{-1}(\gamma_{1}-\gamma_{2})\right)^{\top}\left(\begin{matrix}0\\ 1\\ 2\end{matrix}\right).$ (15)
To prove nonidentifiability of the rankings, it thus suffices to show that for any $\Theta$, there exists $\gamma_{1},\gamma_{2},\Delta$ such that the score difference (15) is zero at $h=0$, i.e. (14) holds, and the derivative of the score difference (15) is nonzero, i.e.
$\nabla_{h}\left(\left(\Theta+h\Delta\right)^{-1}(\gamma_{1}-\gamma_{2})\right)^{\top}\left(\begin{matrix}0\\ 1\\ 2\end{matrix}\right)\propto\left(\Theta^{-1}(\gamma_{1}-\gamma_{2})\right)^{\top}(\Theta^{-1}\Delta)^{\top}\left(\begin{matrix}0\\ 1\\ 2\end{matrix}\right)\neq 0.$ (16)
@@
[…]
@@
Because $\mathbf{1}^{\top}(u-\bar{u}\mathbf{1})a^{T}=\vec{0}^{\top},$ we have constructed a $\Delta$ such that every column sums to 0. Moreover, because $\Theta^{-1}(\gamma_{1}-\gamma_{2})=\epsilon\alpha$, (16) simplifies to
@@
F.4 Proof for asymptotic consistency of independent model
@@
$\prod_{i=1}^{n}\prod_{j=1}^{J}\prod_{k\neq j}\Big[\sum_{m=1}^{M}\underbrace{\Pr\left(\hat{S}^{(j)}_{ik}|S_{ik}^{*}=m;\theta_{m,k}^{(j)}\right)}_{=\theta_{m,k,\hat{S}^{(j)}_{ik}}^{(j)}}\underbrace{\Pr(S_{ik}^{*}=m)}_{=\pi_{k,m}}\Big].$ (17)
Here we prove that the candidate ranking that maximizes (17) still converges asymptotically to the true ranking as long as the true rank is identifiable, despite the probability model ignoring correlation between judge-assigned scores (and despite the fact that $\theta$ and $\pi$ are not identifiable). This can be viewed as an extension of traditional statistical results that the MLE for parameters that only depend on the first-moment is asymptotically consistent even when we do not correctly model the second-moments or higher. This result is stated more formally below. For simplicity, we present the result for one judge that satisfies Assumption 4. Extensions to multiple judges and to use Assumption 5 rather than 4 are straightforward.
@@
Suppose Assumption 4 holds for one judge. The estimated ranking $\hat{\xi}_{n}$ using scores from the judge by maximizing (17) converges to the true ranking $\xi^{*}$ as $n\rightarrow\infty$ as long as:
@@
$\max_{\gamma_{1},\cdots,\gamma_{K}}\prod_{i=1}^{n}\prod_{k=1}^{K}\Pr\left(\hat{S}_{i,k};\gamma_{k}\right).$ (18)
@@
Because the solution to (18) corresponds to the MLE for a multinomial model, and standard regularity conditions apply, the estimator is asymptotically consistent. Formally, this implies that the estimated parameters $\hat{\gamma}_{k,n}$ satisfy
$\Pr\left(\ell_{n}(\gamma_{1},\cdots,\gamma_{K})-\ell_{n}(\gamma_{1}^{*},\cdots,\gamma_{K}^{*})<0\quad\forall\|\gamma-\gamma^{*}\|\geq a\right)\rightarrow 1$ (19)
@@
Let’s analyze the optimization problem (17) optimized with respect to the parameters $(\theta,\delta,\xi)$ where $\xi$ is the ranking of the candidate, $\delta\in\mathbb{R}^{K,+}$, and $\pi_{k}=\sum_{i=1}^{K}\mathbbm{1}\{i<=\xi_{k}\}\delta_{i}$. Let a function $g$ denote the mapping from a parameter set $(\theta,\delta,\xi)$ describing judge and candidates to the set of marginal parameters $\gamma_{1},\cdots,\gamma_{K}$. Under Assumption 4, there is some true set of parameters $(\theta^{*},\delta^{*},\xi^{*})$ with corresponding marginal parameters $\gamma_{1}^{*},\cdots,\gamma_{K}^{*}$. Then because the set of models considered in (17) is a subset of that in (18), we have that
$\Pr\left(\ell_{n}(g(\theta,\delta,\xi))-\ell_{n}(g(\theta^{*},\delta^{*},\xi^{*}))<0\quad\forall\left\|g(\theta,\delta,\xi)-q(\theta^{*},\delta^{*},\xi^{*})\right\|\geq a\right)$ (20)
$\geq$ $\Pr\left(\ell_{n}(\gamma_{1},\cdots,\gamma_{K})-\ell_{n}(\gamma_{1}^{*},\cdots,\gamma_{K}^{*})<0\quad\forall\|\gamma-\gamma^{*}\|\geq a\right).$ (21)
Because (21) converges to one for any $a>0$, then (20) converges asymptotically to one as well for all $a>0$. Taking $a\rightarrow 0$, we have that
$\sup_{\theta_{n},\delta_{n},\xi_{n}}g(\hat{\theta}_{n},\delta_{n},\xi_{n})-(\gamma_{1}^{*},\cdots,\gamma_{K}^{*})\rightarrow_{p}0.$ (22)
Furthermore, because we assumed that (i) the true ranking $\xi^{*}$ is identifiable from the marginals and (ii) no two candidates have the same exact true score, then the mapping from candidate marginals $(\gamma_{1},\cdots,\gamma_{K})$ to candidate rankings is unique and locally constant at $(\gamma_{1}^{*},\cdots,\gamma_{K}^{*})$. Then by the continuous mapping theorem, we have that the estimated rankings $\hat{\xi}_{n}$ when maximizing (17) converges asymptotically to the true rankings $\xi^{*}$. ∎